OpenStudy (thesammysaurus):

A student is assessing the correlation between the number of hours a plant receives sunlight and the height to which it grows. The table below shows the data: Number of hours of sunlight (x) 0 1 2 3 4 5 6 7 8 9 Height of plant in cm (y) 4 7 10 13 16 19 22 25 28 31 Part A: Is there any correlation between the number of hours of sunlight that a plant receives and the height to which it grows? Justify your answer. (4 points) Part B: Write a function which best fits the data. (3 points) Part C: What does the slope and y-intercept of the plot indicate about the plant? (3 points)

jimthompson5910 (jim_thompson5910):

Notice the x values `0 1 2 3 4 5 6 7 8 9` increase by 1 each time at the same time, the y values `4 7 10 13 16 19 22 25 28 31` are increasing by what each time?

OpenStudy (thesammysaurus):

3

jimthompson5910 (jim_thompson5910):

yes these two facts will help us find the slope slope = (change in y)/(change in x) = (+3)/(+1) = 3 slope = 3

OpenStudy (thesammysaurus):

ok so y= 3x+1?

jimthompson5910 (jim_thompson5910):

the slope is m = 3 the y intercept is b = 4 this is the y value that corresponds to x = 0

jimthompson5910 (jim_thompson5910):

now plug m = 3 and b = 4 into y = mx+b to get your linear equation

OpenStudy (thesammysaurus):

y= 3x+4

jimthompson5910 (jim_thompson5910):

yes the last step you do is replace y with f(x) to get it into function notation form

jimthompson5910 (jim_thompson5910):

that will wrap up part B

OpenStudy (thesammysaurus):

Is part A correct?

jimthompson5910 (jim_thompson5910):

what did you write for part A?

OpenStudy (thesammysaurus):

There is a positive correlation between the number of hours of sunlight that a plant receives and the height to which it grows because as the hours increase, the plant grows taller.

jimthompson5910 (jim_thompson5910):

yes that's correct

OpenStudy (thesammysaurus):

I also wrote Part C) That the plant starts at 4 cm which is the why intercept and increases by 3 cm each time which is the slope

jimthompson5910 (jim_thompson5910):

Since a line fits through ALL of the points perfectly, we have a perfect correlation. Aka, a causation (one variable directly affects the other)

OpenStudy (thesammysaurus):

Y-intercept autocorrect oops

OpenStudy (thesammysaurus):

ok thank u

jimthompson5910 (jim_thompson5910):

`That the plant starts at 4 cm` yes basically with no time in the sun, the plant is 4 cm tall the slope of 3 means the plant increases 3 cm for each hour of sunlight

OpenStudy (thesammysaurus):

thank u!

jimthompson5910 (jim_thompson5910):

no problem